# HandelBot: Real-World Piano Playing via Fast Adaptation of Dexterous Robot Policies

### 12 Mar 2026 v1

@@ II RELATED WORK
Robotic Piano Playing. Early robotic piano systems relied on specialized hardware and handcrafted controllers designed explicitly for keyboard actuation [24, 30, 47, 19, 2, 61, 60, 29]. Recent learning-based approaches have shown dexterous piano playing using general-purpose, non-specialized robotic hardware. These approaches leverage data-driven methods, including large-scale human motion capture [50], video-based kinematic retargeting [40], and reinforcement learning in simulation [52, 58].
@@
Real-World Reinforcement Learning. Real-world reinforcement learning holds the potential of autonomous learning under real-world dynamics and environments. However, it remains challenging due to real-world reward functions, environment resets, safety and more. Prior work has shown that real-world RL from scratch is possible with careful design choices [45, 17, 12, 34]. To address challenges with sample inefficiency and exploration in high-dimensional action spaces, prior works bootstrap with prior demonstrations [16, 63], multi-task or offline data [18, 66], or large pretrained models [54, 35]. In our work, we use an open-loop rollout learned from RL in sim, providing a strong initialization and reducing the exploration space.
@@ ACKNOWLEDGMENT
We would like to thank members of ILIAD for their feedback and support. We thank Satvik Sharma and Jayson Meribe for hardware assistance and support, Marcel Torne Villasevil and Megha Srivastava for project brainstorming and simulator development, and Hung-Chieh Fang for hardware and software advice.

### 15 Mar 2026 v2

@@ II RELATED WORK
Robotic Piano Playing. Early robotic piano systems relied on specialized hardware and handcrafted controllers designed explicitly for keyboard actuation [24, 31, 48, 19, 2, 63, 62, 30, 52]. Recent learning-based approaches have shown dexterous piano playing using general-purpose, non-specialized robotic hardware. These approaches leverage data-driven methods, including large-scale human motion capture [51], video-based kinematic retargeting [41], and reinforcement learning in simulation [54, 60].
@@
Real-World Reinforcement Learning. Real-world reinforcement learning holds the potential of autonomous learning under real-world dynamics and environments. However, it remains challenging due to real-world reward functions, environment resets, safety and more. Prior work has shown that real-world RL from scratch is possible with careful design choices [46, 17, 12, 35]. To address challenges with sample inefficiency and exploration in high-dimensional action spaces, prior works bootstrap with prior demonstrations [16, 65, 29], multi-task or offline data [18, 68], or large pretrained models [56, 36]. In our work, we use an open-loop rollout learned from RL in sim, providing a strong initialization and reducing the exploration space.
@@ ACKNOWLEDGMENT
This work was funded by ONR MURI N00014-25-1-2479, and ONR YIP N00014-22-1-2293, NSF #2218760 and NSF #1941722. We would like to thank members of ILIAD for their feedback and support. We thank Satvik Sharma and Jayson Meribe for hardware assistance and support, Marcel Torne Villasevil and Megha Srivastava for project brainstorming and simulator development, Hung-Chieh Fang for hardware and software advice, and Abrar Anwar and Hengyuan Hu for RL advice.